a flower bed length is 3m. more than twice the width. the flower bed has an area of 20 meters squared find the dementions.
Answer by nerdybill(7384)

a flower bed length is 3m. more than twice the width. the flower bed has an area of 20 meters squared find the dementions.
.
Let w = width
then
2w+3 = length
.
w(2w+3) = 20
2w^2+3w = 20
2w^2+3w-20 = 0
2w^2+8w-5w-20 = 0
(2w^2+8w)-(5w+20) = 0
2w(w+4)-5(w+4) = 0
(w+4)(2w-5) = 0
w = 5/2 = 2.5 meters (width)
.
length
2w+3 = 2(5/2)+3 = 5+3 = 8 meters (length)
.
Dimensions are 2.5 meters by 8 meters